OpenStudy (katielong):

write down the equation of a line perpendicular to y=5x+3? I put y=-5x+3 but the teacher said its wrong Thanks

sam (.sam.):

Of course its wrong, :) the gradient is 5,so to find the perpendicular line use the formula \[m_{1}+m_{2}=-1\] 5m=-1 m=-1/5

sam (.sam.):

the equation should be \[m_{1}m_{2}=−1\] not \[m_{1}+m_{2}=−1\]

sam (.sam.):

once you have ,m=-1/5, apply to any coordinate (since you didnt specify any coordinate) At (8,0) \[y-y_{1}=m(x-x_{1})\] y=-1/5 x + 1/5 +8 y=-1/5 x +41/5

OpenStudy (callisto):

At (8,0) \[y-y_{1}=m(x-x_{1})\] y-0 = -1/5 (x-8) y = -1/5x + 8/5

sam (.sam.):

whoops made a mistake there y−y1=m(x−x1) y=-1/5 x + 8/5

OpenStudy (katielong):

no, we haven't done anything like that before... i dont think that would be it

OpenStudy (anonymous):

The teacher is right. Your doing negative reciprocal with perpendicular problems. (For example, the negative reciprocal of 4 is -1/4) You the negative reciprocal for slope only.